OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = x +

OpenStudy (anonymous):

@Godlovesme

Directrix (directrix):

Use the online calculator from before. Let me find it.

OpenStudy (anonymous):

where is it

Directrix (directrix):

y - y1 = m*(x - x1) where m is the slope and (x1, y1) is the given point.

OpenStudy (godlovesme):

i'll let directix help u

Directrix (directrix):

y - 5 = -4*(x - ( -1))

OpenStudy (anonymous):

so y=-4x- (-4)

Directrix (directrix):

Just a sec. Check the signs.

OpenStudy (anonymous):

y=4x+(-1)

Directrix (directrix):

The slope is - 4.